OpenStudy (anonymous):

find the derivative of the function. f(t)= (3+6t)^(3/7)

OpenStudy (anonymous):

You know chain rule and product rule?

OpenStudy (anonymous):

i cant do the simple power rule? cause i got (3/7)(3+6t)^(-4/7) so far but idk where to go from here and i know this is wrong

OpenStudy (anonymous):

\[f(t)= (3+6t)^{3/7}\] \[f \prime (t)= \frac{ 3 }{ 7 }(3+6t)^{\frac{ 3 }{ 7 } - 1} \times 6\]

OpenStudy (anonymous):

you get it?

OpenStudy (anonymous):

what rule is that?

OpenStudy (anonymous):

this is called the chain rule

OpenStudy (anonymous):

i dont get why you multiply 6 can you explain it to me please

OpenStudy (anonymous):

uhm do you know he simple x^n differentiation?

OpenStudy (anonymous):

yeah the simple power rule i did that but it gave me the wrong answer

OpenStudy (anonymous):

ya now you assume the ( 3 + 6t ) = x and then differentiate it

OpenStudy (anonymous):

gottcha okay i understand now so basically you do the simple power rule and then multiply what is outside of the parenthesis by the derivative of inside the parenthesis

OpenStudy (anonymous):

(Sorry, when I said "chain rule and product rule," I meant chain rule and power rule.)

OpenStudy (anonymous):

@CliffSedge thats okay thanks for the help!

OpenStudy (anonymous):

@miteshchvm ??

OpenStudy (anonymous):

oh ha ha! you understood it?

OpenStudy (anonymous):

"gottcha okay i understand now so basically you do the simple power rule and then multiply what is outside of the parenthesis by the derivative of inside the parenthesis" so this is right?? @miteshchvm haha ? :D

OpenStudy (anonymous):

ya absolutely right! :D

OpenStudy (anonymous):

okay thanks a bunch for your help!!! <3 :DDDD